Assume that AD is the geometric mean of BD and DC in triangle ABC.
a) Find AD if BD = 6 and DC = 8
b) Find BD if AD = 6 and DC = 8

Answer by stanbon(75887)

Assume that AD is the geometric mean of BD and DC in triangle ABC.
BD/AD = AD/DC
------
a) Find AD if BD = 6 and DC = 8
6/AD = AD/8
AD = sqrt(48) = 4sqrt(3)
=================================
b) Find BD if AD = 6 and DC = 8
BD/6 = 6/8
BD = 36/8
BD = 9/2
---------------------
Cheers,
Stan H.
---------------------